OpenStudy (anonymous):

can someone help I need help with making fractions or mixed numbers into decimals

OpenStudy (anonymous):

you can divide the fraction with "long division" and find the decimal value that way

OpenStudy (anonymous):

umm is that what im supposed to do our just asking

OpenStudy (anonymous):

yes take example fraction

OpenStudy (anonymous):

hold on really quickly I have to shut down the coputer just for a minute because the email sign is not working

OpenStudy (anonymous):

for me

OpenStudy (debbieg):

It kind of depends on the fraction, also. If you can get the den'r to a power of 10, then you can easily convert to a decimal. E.g.: convert 4/5 to decimal. \[\Large \dfrac{ 4 }{ 5 }=\dfrac{ 4 }{ 5 }\cdot \dfrac{ 2 }{ 2 }=\dfrac{ 8 }{ 10 }=0.8\]

OpenStudy (anonymous):

thankyou is that how I write he problem and thankyou again

OpenStudy (debbieg):

Well, if the fraction you are to convert is 4/5, yeah, that's how I'd write it... lol. Like I said, that only works if the den'r is a factor of a power of 10. It won't work with every fraction. Like @phanta_seea said, sometimes you may need to do long division. That will always work.

OpenStudy (debbieg):